Dawn is m years old now. She has two older brothers, Eugene and Aaron, who often look after her. Eugene is 6 times as old as Dawn. Aaron is 11 years older than Eugene. How old was Aaron 5 years ago?
Found 2 solutions by Theo, josgarithmetic:
Answer by Theo(13342)

dawn is m year old now.

eugene is 6 times as old as dawn.
that makes eugene 6m years old.

aaron is 11 years older than eugene.
that makes aaron 6m + 11 years old.

how old was aaron 5 years ago?
5 years ago, aaron was 6m + 11 - 5 = 6m + 6 years old.
that's your solution.

for example:
assume dawn is 2 years old now.
that makes eugene 6 * 2 = 12 years old now.
that makes aaron 12 + 11 = 23 years old now.
5 years ago, aaron was 23 - 5 = 18 years old.
6 * 2 + 6 = 12 + 6 = 18, confirming the solution is correct.


Answer by josgarithmetic(39620)

PERSON           AGE NOW
Eugene             6m
Aaron              6m+11
Dawn                m

Aaron five years ago was 6m+6.